OpenStudy (anonymous):

Simplify: Simplify: 64/169 A. 13/8 B. 8/13 C. 32/84 D. 64/169 A. 13/8 B. 8/13 C. 32/84 D. 64/169 i was just gottna put d im not sure help!!

OpenStudy (anonymous):

\[\sqrt{64/169}\]

OpenStudy (anonymous):

sqrt(64/169) =sqrt((8/13)^2) =8/13